Documenta Mathematica

Band 7, 2002

Kengo Matsumoto C^{∗}-Algebras Associated

with Presentations of Subshifts  1–30

Gordon Heier

Effective Freeness of Adjoint Line Bundles  31–42

Ezra Miller Planar Graphs as Minimal Resolutions of

Trivariate Monomial Ideals  43–90

J. Brodzki, R. Plymen

Complex Structure on the Smooth Dual of GL(n)  91–112

Victor Reiner

Equivariant Fiber Polytopes  113–132

Shoji Yokura On the Uniqueness Problem

of Bivariant Chern Classes  133–142

B. Mirzaii, W. van der Kallen

Homology Stability for Unitary Groups  143–166

Raymond Brummelhuis, Heinz Siedentop, and Edgardo Stockmeyer The Ground State Energy of Relativistic One-Electron Atoms

According to Jansen and Heß  167–182

Frans Oort and Thomas Zink Families of p-Divisible Groups

with Constant Newton Polygon  183–201

Paul Balmer, Stefan Gille, Ivan Panin and Charles Walter The Gersten Conjecture for Witt Groups

in the Equicharacteristic Case  203–217

Martin Olbrich

L^{2}-Invariants of Locally Symmetric Spaces  219–237

Cornelia Busch

The Farrell Cohomology of SP(p − 1, ℤ)  239–254

Guihua Gong On the Classification of Simple Inductive Limit C^{∗}-Algebras, I:

The Reduction Theorem  255–461

Winfried Bruns and Joseph Gubeladze

Unimodular Covers of Multiples of Polytopes  463–480

N. Karpenko and A. Merkurjev

Rost Projectors and Steenrod Operations  481–493

Tamás Hausel and Bernd Sturmfels

Toric Hyperkähler Varieties  495–534

Bernhard Keller and Amnon Neeman The Connection Between May’s Axioms for a Triangulated Tensor Product and Happel’s Description

of the Derived Category of the Quiver D_{4}  535–560

Anthony D. Blaom Reconstruction Phases for

Hamiltonian Systems on Cotangent Bundles  561–604

Marc A. Rieffel Group C^{∗}-Algebras as Compact

Quantum Metric Spaces  605–651

Georg Schumacher Asymptotics of Complete Kähler-Einstein Metrics – Negativity of the

Holomorphic Sectional Curvature  653–658